# HiVG: Hierarchical Multimodal Fine-grained Modulation for Visual Grounding

### v1

@@ Abstract.
Visual grounding, which aims to ground a visual region via natural language, is a task that heavily relies on cross-modal alignment. Existing works utilized uni-modal pre-trained models to transfer visual/linguistic knowledge separately while ignoring the multimodal corresponding information. Motivated by recent advancements in contrastive language-image pre-training and low-rank adaptation (LoRA) methods, we aim to solve the grounding task based on multimodal pre-training. However, there exists significant task gaps between pre-training and grounding. Therefore, to address these gaps, we propose a concise and efficient hierarchical multimodal fine-grained modulation framework, namely HiVG. Specifically, HiVG consists of a multi-layer adaptive cross-modal bridge and a hierarchical multimodal low-rank adaptation (Hi LoRA) paradigm. The cross-modal bridge can address the inconsistency between visual features and those required for grounding, and establish a connection between multi-level visual and text features. Hi LoRA prevents the accumulation of perceptual errors by adapting the cross-modal features from shallow to deep layers in a hierarchical manner. Experimental results on five datasets demonstrate the effectiveness of our approach and showcase the significant grounding capabilities as well as promising energy efficiency advantages. The project page: https://github.com/linhuixiao/HiVG.
@@
copyright: acmlicensedjournalyear: 2024conference: Proceedings of the ACM International Conference on Multimedia; 2024; ArXivisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Visual Grounding (VG), also known as Referring Expression Comprehension (REC) or Phrase Grounding (Qiao et al., 2020; Mao et al., 2016; Yu et al., 2016; Hu et al., 2016; Deng et al., 2021; Xiao et al., 2023, 2019; Li et al., 2019), is a fundamental and challenging task at the intersection fields of vision-language understanding, which can be potentially used in a wide range of applications, such as visual question answering (Antol et al., 2015) and human-machine interaction (Chen et al., 2023b) etc.. Unlike object detection (Liu et al., 2023b, c), which requires a predefined and fixed set of categories, grounding is not limited to specific categories but instead needs to identify the specific image region according to the language expression semantics. Thus, grounding is a task that strongly relies on the interaction and alignment of multimodal features.
@@
By utilizing language supervision from large-scale unlabeled data, Vision-Language Pre-training (VLP) can acquire comprehensive multimodal representations. Recently, the remarkable success of Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) has demonstrated its ability to learn general visual concepts, which assists many multimodal tasks to achieve remarkable improvements (Radford et al., 2021; Peng et al., 2023b; Tao et al., 2023; Kim et al., 2023; Wang et al., 2022a). In visual grounding, there are also works, e.g., CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), which consider using CLIP. However, existing methods mainly utilize the CLIP as a backbone to extract strong vision and language features, without comprehensively investigating on the significant task gap between the pre-trained CLIP and the downstream grounding, which hinders exploiting the full potential of pre-training models. In this work, we scrutinize the task gap from two aspects. (1) Data bias. There inevitably exists a certain bias in data between the large-scale pre-training and grounding. Directly utilizing the frozen vision backbone of the CLIP may extract visual features sensitive to general objects that are not the focus of the query in visual grounding. For example, as shown in Fig. 1-(a), the middle giraffe receives highlight attention, but it has little relation to the grounding task. (2) Difference in learning objectives. The visual grounding task needs to find the precise image region that has the target object expressed by the query sentence. In contrast, CLIP works as a multimodal pre-trained model, which is only constrained to coarsely align noisy image and text data (Schuhmann et al., 2021) in a self-supervised way. In addition, the self-supervised constraint is only performed at the final layer. When directly using the pre-trained CLIP in visual grounding, some valuable fine-grained visual information in the bottom vision layers may be discarded, which brings challenges for accurately locating the object box. For example, as shown in Fig. 1-(a), the left giraffe receives relatively small attention areas, which leads to inaccurate box of the target object.
@@
In this paper, we propose a hierarchical multimodal fine-grained modulation framework to more effectively adapt the pre-trained CLIP to grounding, namely HiVG. It is a concise and efficient end-to-end framework that can simultaneously alleviate two kinds of task gaps (i.e., data bias and learning objectives) through a multi-layer adaptive cross-modal bridge and a hierarchical low-rank adaptation paradigm.
Firstly, to address the inconsistency between visual features of the pre-trained CLIP and those required for grounding, as well as establish a connection between multi-level visual and text features, we have designed a multi-layer adaptive cross-modal bridge. Specifically, the cross-modal bridge includes a sample-agnostic semantic weighting module and a multi-head cross-attention module. The weighting module incorporates learnable multi-level sample-agnostic adaptive weights, facilitating the selection of appropriate linguistic features through a residual operation. The multi-head cross-attention utilizes the selected multi-level text features for guiding the learning of the visual features required in grounding. The sample-agnostic semantic weighting scheme is inspired by (Cao et al., 2020), i.e., specific layers of a pre-trained model may have distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers.
@@ 2.1. Visual Grounding
Visual grounding has recently received significant research attention, and it can be categorized into several settings. On the one hand, represented by TransVG (Deng et al., 2021), this setting involves full-parameter fine-tuning utilizing pre-trained closed-set detectors and language models. It is considered the most conventional and extensively studied setting. Under this setting, numerous complex two-stage (Yu et al., 2018; Liu et al., 2019b, c; Hong et al., 2019) and one-stage (Yang et al., 2020, 2019a; Zhou et al., 2021) methods emerged based on traditional detection networks in the early CNN era. After the introduction of ViT (Vaswani et al., 2017; Dosovitskiy et al., 2020), the Transformer-based networks (Deng et al., 2021; Ye et al., 2022a; Kamath et al., 2021; Ho et al., 2023; Li and Sigal, 2021; Xiong et al., 2023; Su et al., 2023; Deng et al., 2023; Ye et al., 2022b) constantly pushes the accuracy to new limits. However, these works only focus on achieving grounding by using independently pre-trained uni-modal detectors and language encoders while ignoring the alignment of cross-modality information within pre-trained model itself. More recent works, such as QRNet (Ye et al., 2022a), VG-LAW (Su et al., 2023), TransVG++ (Deng et al., 2023), etc., only incorporate language-guided knowledge in vision backbone without attempting multi-level fine-grained alignment of multimodal features. Motivated by this setting, several works, such as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), have recently sprung up to the setting of fine-tuning with vision and language self-supervised pre-trained models. Following this setting, our study delves into a deeper perspective of hierarchical multimodal information and achieves fine-grained interaction of cross-modal features. On the other hand, with the evolution of the pre-training paradigm, many new settings have recently emerged that significantly improve the grounding performance, such as fine-tuning with box-level dataset-mixed open-set detection pre-trained models (e.g., MDETR (Kamath et al., 2021), Grounding-DINO (Liu et al., 2023d), etc.), fine-tuning with box-level / multi-task mixup-supervised pre-trained models (e.g., UniTAB (Yang et al., 2022), UNITER (Chen et al., 2020), OFA (Wang et al., 2022b), ONE-PEACE(Wang et al., 2023b), etc.), and grounding multimodal large language models (GMLLM, e.g., Shikra (Chen et al., 2023b), Kosmos-2 (Peng et al., 2023a), Ferret (You et al., 2024), LION (Chen et al., 2023a), etc.). However, these works require a large amount of fine-grained labeled data, resulting in a relatively high training cost.
@@ 2.2. Contrastive Language-Image Pre-training
With the promotion of learning general and transferable cross-modal representations, VLP has become the core training paradigm and research foundation of modern Vision-Language (V-L) research. Benefiting from self-supervised contrastive learning, CLIP has demonstrated impressive generalization and downstream transfer ability in a series of studies (Peng et al., 2023b; Tao et al., 2023). More recently, some works utilized CLIP to realize grounding transfer, such as adapting-CLIP (Li et al., 2022b), ReCLIP (Subramanian et al., 2022) etc., but these works are limited to using CLIP features as aids in an unsupervised or zero-shot setting and cannot directly perform grounding. Although CLIP-VG (Xiao et al., 2023), Dynamic-MDETR (Shi et al., 2022), JMR (Zhu et al., 2023) etc., realizes grounding transfer, it does not conduct more in-depth research on the task gaps and the hierarchical cross-modal features. Unlike previous work, our study fills the gap by conducting a more comprehensive study of the cross-modal task gaps between CLIP’s pre-training and downstream grounding.
@@ 2.3. Low-Rank Adaptation
LoRA (Hu et al., 2021) freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture, thereby significantly reducing the number of trainable parameters for downstream tasks. Vanilla LoRA has been proposed in the field of natural language processing for Large Language Models (LLM) such as LLaMA2 (Touvron et al., 2023), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020) with 175B parameters, etc.. Recently, researchers have attempted to apply vanilla LoRA in the fields of cross-modal tasks (Smith et al., 2023). However, since cross-modal tasks primarily emphasize the interaction of multimodal information in contrast to unimodal language or visual tasks, the application of LoRA to grounding tasks remains unexplored. Consequently, we propose Hi LoRA as an effective solution for addressing the existing gaps in multimodal downstream transfer.
@@ 3.2. Multi-layer Adaptive Cross-modal Bridge
The visual encoder of CLIP independently encodes the image, and the obtained multi-level visual features may be inconsistent with those required for grounding. Additionally, as inspired by (Cao et al., 2020), specific layers of a pre-trained model may exhibit distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers. Therefore, we should provide a wide range of multi-level text features for different visual layers to select and calibrate. Thus, to address these issues, we propose integrating a multi-layer adaptive cross-modal bridge into the image encoder to achieve fine-grained visual features.
@@ 3.3. Hierarchical Low-Rank Adaptation
HiVG (ours) CLIP-B CLIP-B 87.32 89.86 83.27 78.06 84.81 68.11 78.29 78.79 75.22 82.11
HiVG-L (ours) CLIP-L CLIP-L 88.14 91.09 83.71 80.10 86.77 70.53 80.78 80.25 76.23 82.16
@@
UniTAB (Yang et al., 2022) ECCV’22 RN101+DETR RoBERT-B 86.32 88.84 80.61 78.70 83.22 69.48 79.96 79.97 79.38
@@
HiVG (ours) CLIP-B CLIP-B 90.56 92.55 87.23 83.08 89.21 76.68 86.52 86.62 77.75 82.08
HiVG-L (ours) CLIP-L CLIP-L 91.37 93.64 88.03 86.78 89.91 78.02 88.61 88.60 78.16 82.63
@@ 4.1. Implementation Details
Datasets and Evaluation Metrics. The effectiveness of our method is validated on five widely utilized datasets, namely the three REC datasets (RefCOCO/+/g (Yu et al., 2016; Mao et al., 2016)), as well as two Phrase Grounding (PG) datasets (ReferItGame (Kazemzadeh et al., 2014a) and Flickr30k Entities (Plummer et al., 2015)). In PG, the query pertains to a specific phrase, while in REC, the query refers to a reference expression. The text of RefCOCO+/g exhibits greater length and complexity in comparison to that of RefCOCO. We follow the previous research that employs Intersection-over-Union (IoU) as the evaluation metric. Specifically, a prediction is deemed accurate only when its IoU exceeds or equals 0.5. Finally, we compute the prediction accuracy for each dataset as a performance indicator. The detailed description and statistics information regarding these five datasets are provided in the Appendix.
Network Architecture. We employed CLIP ViT-B/16 and CLIP ViT-L/14 as the backbone of our HiVG-B and HiVG-L versions. In the based version, the Hi LoRA module utilizes a rank of 32 and an $\alpha$ coefficient of 16. The encoder layers are evenly divided into 3 groups, and Hi LoRA is applied with 3 stages accordingly. HiVG extracted 1th, 4th, 8th, and 12th layer features of the visual encoder, the cross-modal bridge injected 4th, 8th, and 12th layer, and text aggregated from 1th to 12th layer features of the text encoder. In the grounding encoder, we adopted the pre-norm instead of the post-norm structure and set the hidden dimensions as the same with text encoder. The detailed structural parameters of the framework and the large version are summarized in the Appendix.
@@
Training Details. To prevent catastrophic forgetting, we freeze the original parameters of CLIP’s visual and textual encoders. Since the parameters of the low-stage Hi LoRA are included in the high-stage Hi LoRA, our updated parameters do not show any increase compared to the vanilla LoRA. Besides, Hi LoRA represents a PEFT approach for the pre-trained model, and the grounding encoder employs random Xavier initialization. Thus, to enhance training stability, we perform training in two stages. In the first stage, we trained the grounding encoder, regression head at a high learning rate without activating Hi LoRA. It is imperative to employ Hi LoRA for the text encoder with only one layer group as well, in order to mitigate the risk of catastrophic forgetting. The batch size is set to 80. Our model is optimized end-to-end by using the AdamW optimizer and a cosine learning scheduler with an initial learning rate of $2.5\times 10^{-4}$ for 50 epochs during the first stage. During Hi LoRA adaptation, the learning rates in three stages are $1.0\times 10^{-4}$, $0.5\times 10^{-4}$, and $0.25\times 10^{-4}$ with 20 epochs, respectively. Our framework and experiments are based on PyTorch by using 8 NVIDIA A100 GPUs. More details are provided in the Appendix.

### v2

@@ Abstract.
Visual grounding, which aims to ground a visual region via natural language, is a task that heavily relies on cross-modal alignment. Existing works utilized uni-modal pre-trained models to transfer visual or linguistic knowledge separately while ignoring the multimodal corresponding information. Motivated by recent advancements in contrastive language-image pre-training and low-rank adaptation (LoRA) methods, we aim to solve the grounding task based on multimodal pre-training. However, there exists significant task gaps between pre-training and grounding. Therefore, to address these gaps, we propose a concise and efficient hierarchical multimodal fine-grained modulation framework, namely HiVG. Specifically, HiVG consists of a multi-layer adaptive cross-modal bridge and a hierarchical multimodal low-rank adaptation (HiLoRA) paradigm. The cross-modal bridge can address the inconsistency between visual features and those required for grounding, and establish a connection between multi-level visual and text features. HiLoRA prevents the accumulation of perceptual errors by adapting the cross-modal features from shallow to deep layers in a hierarchical manner. Experimental results on five datasets demonstrate the effectiveness of our approach and showcase the significant grounding capabilities as well as promising energy efficiency advantages. The project page: https://github.com/linhuixiao/HiVG.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681071isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Scene understanding
@@ 1. Introduction
Visual Grounding (VG), also known as Referring Expression Comprehension (REC) or Phrase Grounding (PG) (Qiao et al., 2020; Mao et al., 2016; Yu et al., 2016; Hu et al., 2016; Deng et al., 2021; Xiao et al., 2023, 2019; Li et al., 2019), is a fundamental and challenging task at the intersection fields of vision-language understanding, which can be potentially used in a wide range of applications (Liu et al., 2023b; Antol et al., 2015; Chen et al., 2023b), such as visual question answering (Antol et al., 2015), human-machine interaction (Chen et al., 2023b) etc.. Unlike object detection (Liu et al., 2023c, d), which requires a predefined and fixed set of categories, grounding is not limited to specific categories but instead needs to identify the specific image region according to the language expression semantics. Thus, grounding is a task that strongly relies on the interaction and alignment of multimodal features.
@@
By utilizing language supervision from large-scale unlabeled data, Vision-Language Pre-training (VLP) can acquire comprehensive multimodal representations. Recently, the remarkable success of Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) has demonstrated its ability to learn general visual concepts, which assists many multimodal tasks to achieve remarkable improvements (Radford et al., 2021; Peng et al., 2023b; Kim et al., 2024; Wang et al., 2022a). In visual grounding, there are also works, e.g., CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), which consider using CLIP. However, existing methods mainly utilize the CLIP as a backbone to extract strong vision and language features, without comprehensively investigating on the significant task gap between the pre-trained CLIP and the downstream grounding, which hinders exploiting the full potential of pre-training models. In this work, we scrutinize the task gap from two aspects. (1) Data bias. There inevitably exists a certain bias in data between the large-scale pre-training and grounding. Directly utilizing the frozen vision backbone of the CLIP may extract visual features sensitive to general objects that are not the focus of the query in visual grounding. For example, as shown in Fig. 1-(a), the middle giraffe receives highlight attention, but it has little relation to the grounding task. (2) Difference in learning objectives. The visual grounding task needs to find the precise image region that has the target object expressed by the query sentence. In contrast, CLIP works as a multimodal pre-trained model, which is only constrained to coarsely align noisy image and text data (Schuhmann et al., 2021) in a self-supervised way. In addition, the self-supervised constraint is only performed at the final layer. When directly using the pre-trained CLIP in visual grounding, some valuable fine-grained visual information in the bottom vision layers may be discarded, which brings challenges for accurately locating the object box. For example, as shown in Fig. 1-(a), the left giraffe receives relatively small attention areas, which leads to inaccurate box of the target object.
@@
In this paper, we propose a hierarchical multimodal fine-grained modulation framework to more effectively adapt the pre-trained CLIP to grounding, namely HiVG. It is a concise and efficient end-to-end framework that can alleviate two kinds of task gaps (i.e., data bias and learning objectives) through a multi-layer adaptive cross-modal bridge and a hierarchical low-rank adaptation paradigm.
Firstly, to address the inconsistency between visual features of the pre-trained CLIP and those required for grounding, as well as establish a connection between multi-level visual and text features, we have designed a multi-layer adaptive cross-modal bridge. Specifically, the cross-modal bridge includes a sample-agnostic semantic weighting module and a multi-head cross-attention module. The weighting module incorporates learnable multi-level sample-agnostic adaptive weights, facilitating the selection of appropriate linguistic features through a residual operation. The multi-head cross-attention utilizes the selected multi-level text features for guiding the learning of the visual features required in grounding. The sample-agnostic semantic weighting scheme is inspired by (Dar et al., 2023; Cao et al., 2020), i.e., specific layers of a pre-trained model may have distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers.
@@ 2.1. Visual Grounding
Visual grounding has recently received significant research attention, and it can be categorized into several settings. On the one hand, represented by TransVG (Deng et al., 2021), this setting involves full-parameter fine-tuning utilizing pre-trained closed-set detectors and language models. It is considered the most conventional and extensively studied setting. Under this setting, numerous complex two-stage (Yu et al., 2018; Liu et al., 2019b, c; Hong et al., 2019) and one-stage (Yang et al., 2020, 2019a; Zhou et al., 2021) methods emerged based on traditional detection networks in the early CNN era. After the introduction of ViT (Vaswani et al., 2017a; Dosovitskiy et al., 2020), the Transformer-based networks (Deng et al., 2021; Ye et al., 2022a; Kamath et al., 2021; Ho et al., 2023; Li and Sigal, 2021; Su et al., 2023; Deng et al., 2023; Ye et al., 2022b; Lu et al., 2024; Miao et al., 2023) constantly pushes the accuracy to new limits. However, these works only focus on achieving grounding by using independently pre-trained uni-modal detectors and language encoders while ignoring the alignment of cross-modality information within pre-trained model itself. More recent works, such as QRNet (Ye et al., 2022a), VG-LAW (Su et al., 2023), TransVG++ (Deng et al., 2023), etc., only incorporate language-guided knowledge in vision backbone without attempting multi-level fine-grained alignment of multimodal features. Motivated by this setting, several works, such as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), have recently sprung up to the setting of fine-tuning with vision and language (VL) self-supervised pre-trained models. Following this setting, our study delves into a deeper perspective of hierarchical multimodal information and achieves fine-grained interaction of cross-modal features. On the other hand, with the evolution of the pre-training paradigm, many new settings have recently emerged that significantly improve the grounding performance, such as fine-tuning with box-level dataset-mixed open-set detection pre-trained models (e.g., MDETR (Kamath et al., 2021), Grounding-DINO (Liu et al., 2023e), etc.), fine-tuning with box-level / multi-task mixup-supervised pre-trained models (e.g., UniTAB (Yang et al., 2022a), UNITER (Chen et al., 2020), OFA (Wang et al., 2022b), etc.), and grounding multimodal large language models (GMLLMs, e.g., Shikra (Chen et al., 2023b), Kosmos-2 (Peng et al., 2023a), Ferret (You et al., 2024), LION (Chen et al., 2023a), etc.). However, these works require a large amount of fine-grained labeled data, resulting in a relatively high training cost.
@@ 2.2. Contrastive Language-Image Pre-training
With the promotion of learning general and transferable cross-modal representations (Xiong et al., 2024; Yang et al., 2022b, 2021; Guo et al., 2024; Xiong et al., 2023; Li et al., 2024), VLP has become the core training paradigm of modern VL research. Benefiting from self-supervised contrastive learning, CLIP has demonstrated impressive generalization and downstream transfer ability in a series of studies (Peng et al., 2023b; Radford et al., 2021). More recently, some works utilized CLIP to realize grounding transfer, such as adapting-CLIP (Li et al., 2022b), ReCLIP (Subramanian et al., 2022) etc., but these works are limited to using CLIP features as aids in an unsupervised or zero-shot setting (Subramanian et al., 2022; Ke et al., 2023) and cannot directly perform grounding. Although CLIP-VG (Xiao et al., 2023), Dynamic-MDETR (Shi et al., 2022), JMR (Zhu et al., 2023) etc., realizes grounding transfer, it does not conduct more in-depth research on the task gaps and the hierarchical cross-modal features. Unlike previous work, our study fills the gap by conducting a more comprehensive study of the cross-modal task gaps between CLIP’s pre-training and downstream grounding.
@@ 2.3. Low-Rank Adaptation
LoRA (Hu et al., 2021) freezes the weights of pre-trained model and injects trainable rank decomposition matrices into each layer of the Transformer (Vaswani et al., 2017b), thereby significantly reducing the number of trainable parameters for downstream tasks. Vanilla LoRA has been proposed in the field of natural language processing for Large Language Models (LLM) such as LLaMA2 (Touvron et al., 2023), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020) with 175B parameters, etc.. Recently, researchers have attempted to apply vanilla LoRA in the fields of cross-modal tasks (Smith et al., 2023). However, since cross-modal tasks primarily emphasize the interaction of multimodal information in contrast to unimodal language or visual tasks, the application of LoRA to grounding tasks remains unexplored. Consequently, we propose HiLoRA as an effective solution for addressing the existing gaps in multimodal downstream transfer.
@@ 3.2. Multi-layer Adaptive Cross-modal Bridge
The visual encoder of CLIP independently encodes the image, and the obtained multi-level visual features may be inconsistent with those required for grounding. Additionally, as inspired by (Dar et al., 2023; Cao et al., 2020), specific layers of a pre-trained model may exhibit distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers. Therefore, we should provide a wide range of multi-level text features for different visual layers to select and calibrate. Thus, to address these issues, we propose integrating a multi-layer adaptive cross-modal bridge into the image encoder to achieve fine-grained visual features.
@@ 3.3. Hierarchical Low-Rank Adaptation
HiVG (ours) ACM MM’24 CLIP-B CLIP-B 87.32 89.86 83.27 78.06 83.81 68.11 78.29 78.79 75.22 82.11
HiVG-L (ours) ACM MM’24 CLIP-L CLIP-L 88.14 91.09 83.71 80.10 86.77 70.53 80.78 80.25 76.23 82.16
@@
UniTAB (Yang et al., 2022a) ECCV’22 RN101+DETR RoBERT-B 86.32 88.84 80.61 78.70 83.22 69.48 79.96 79.97 79.38
@@
HiVG (ours) ACM MM’24 CLIP-B CLIP-B 90.56 92.55 87.23 83.08 89.21 76.68 84.52 85.62 77.75 82.08
HiVG-L (ours) ACM MM’24 CLIP-L CLIP-L 90.77 92.94 88.03 86.78 89.91 78.02 86.61 86.60 78.16 82.63
@@ 4.1. Implementation Details
Datasets and Evaluation Metrics. The effectiveness of our method is validated on five widely utilized datasets, namely the three REC datasets (RefCOCO/+/g (Yu et al., 2016; Mao et al., 2016)), as well as two PG datasets (ReferItGame (Kazemzadeh et al., 2014a) and Flickr30k Entities (Plummer et al., 2015)). In PG, the query pertains to a specific phrase, while in REC, the query refers to a referring expression. The text of RefCOCO+/g exhibits greater length and complexity in comparison to that of RefCOCO. We follow the previous researches that employs Intersection-over-Union (IoU) as the evaluation metric. Specifically, a prediction is deemed accurate only when its IoU exceeds or equals 0.5. Finally, we compute the prediction accuracy for each dataset as a performance indicator.
Network Architecture. We employed CLIP ViT-B/16 and CLIP ViT-L/14 as the backbone of our HiVG-B (default) and HiVG-L versions. In the base version, the HiLoRA module utilizes a rank of 32 and an $\alpha$ coefficient of 16. The encoder layers are evenly divided into 3 groups, and HiLoRA is applied with 3 stages accordingly. HiVG extracted 1th, 4th, 8th, and 12th layer features of the visual encoder, the cross-modal bridge injected 4th, 8th, and 12th layer, and text aggregated from 1th to 12th layer features of the text encoder. In the grounding encoder, we adopted the pre-norm instead of the post-norm structure and set the hidden dimensions as the same with text encoder.
@@
Training Details. To prevent catastrophic forgetting, we freeze the original parameters of CLIP’s two encoders. Since the parameters of the low-stage HiLoRA are included in the high-stage HiLoRA, our updated parameters do not show any increase compared to the vanilla LoRA. Besides, HiLoRA represents a PEFT approach for the pre-trained model, and the grounding encoder employs random Xavier initialization. Thus, to enhance training stability, we perform training in two stages. In the first stage, we trained the grounding encoder, regression head at a high learning rate without activating HiLoRA. It is imperative to employ HiLoRA for the text encoder with only one layer group as well, in order to mitigate the risk of catastrophic forgetting. The batch size is set to 60. Our model is optimized end-to-end by using the AdamW optimizer and a cosine learning scheduler with an initial learning rate of $2.5\times 10^{-4}$ for 50 epochs during the first stage. During HiLoRA adaptation, the learning rates in three stages are $1.0\times 10^{-4}$, $0.5\times 10^{-4}$, and $0.25\times 10^{-4}$ with 20 epochs, respectively. Besides, to ensure a fair comparison, like the existing works (Deng et al., 2023; Su et al., 2023), we pre-perform a vanilla LoRA adapting of CLIP’s image encoder under ViT-Det (Li et al., 2022a) detection framework on MSCOCO dataset, with excluding the validation and test images of RefCOCO/+/g. Our framework and experiments are based on PyTorch by using 8 NVIDIA A100 GPUs.